The cost of water(in dollars) is a linear fuction of the amount of water used(in hundreds of cubic feet,HCF). The cost of using 18 HCF is $39.35 and the cost of using 55 HCF is $104.10. What is the cost of using 53 HCF of water
Answer by stanbon(75887)

The cost of water(in dollars) is a linear fuction of the amount of water used(in hundreds of cubic feet,HCF). The cost of using 18 HCF is $39.35 and the cost of using 55 HCF is $104.10. What is the cost of using 53 HCF of water
-----------------------
You have two points relating usage and cost.
You want cost to be the dependent variable.
-----
Points: (18 , 39.35) and (55 , 104.10)
-----
slope = (104.10-39.35)/(55-18) = 64.15/37 = 1.7338
intercept: 39.35 = 1.7338*18 + b
b = 8.14189
------------
Equation: cost = 1..7338(use)+8.14189
------------------------------------------
cost(53) = 1.7338*53+8.14189 = $100.03
------------------------------------------
Cheers,
Stan H.